OpenStudy (anonymous):

Use the Remainder Theorem and synthetic division to find P(r). P(x)=2x^4-3x^3+6, r=2

OpenStudy (anonymous):

using remainder theorem P(2)=2(2)^4-3(2)^3+6=32-24+6=14

OpenStudy (anonymous):

by synthetic division 2) 2 -3 0 0 6 0 4 2 4 8 ---------------- 2 1 2 4 14 14 is the remainder

OpenStudy (anonymous):

i dont know u understand it or not :S

OpenStudy (anonymous):

Now that you did it that way, it is much more understanding. The longer, the more confusing